OpenStudy (turingtest):

Suppose that \(f\) satisfies the equation \[f(x+y)=f(x)+f(y)+x^2y+xy^2\]for all real numbers \(x\) and \(y\). Suppose further that\[~~~~~~~~~~~~~\lim_{x\to0}\frac{f(x)}x=1\]a) Find \(f(0)\) b) Find \(f'(0)\) c) Find \(f'(x)\)

OpenStudy (turingtest):

I found \(f(0)=0\) but now I am not seeing how to proceed. Just a hint would be nice.

OpenStudy (turingtest):

I further think I proved that the function is odd, though I think that doesn't really help me :/

OpenStudy (zarkon):

use the definition of a derivative

OpenStudy (zarkon):

the fact that it is odd does seem to help with the last part

OpenStudy (turingtest):

I was thinking of using the definition of the derivative somehow but couldn't figure out how to get it into the right form....

OpenStudy (zarkon):

\[f'(0)=\lim_{x\to0}\frac{f(x)-f(0)}{x-0}\]

OpenStudy (turingtest):

limit f(x)-f(0) something...

OpenStudy (turingtest):

ah dang but... oh, okay I guess I was trying to get things algebraically manipulated there, but I see now that since x=f(x) at x=0 the limit \(is\) the definition got it :)

OpenStudy (turingtest):

okay so that is f'(0)=1, on to the last part you say the oddness could help?

OpenStudy (zarkon):

yes

OpenStudy (turingtest):

can I use the derivative of an odd function is even?

OpenStudy (zarkon):

don't need it

OpenStudy (zarkon):

use the definition of a derivative again

OpenStudy (turingtest):

that's what I'm doing, let me toy for a sec...

OpenStudy (zarkon):

I found f(x) ;)

OpenStudy (turingtest):

haha, dang okay then go ahead and give me a little more help on finding f'(x) they don't even ask for f(x), jeez

OpenStudy (turingtest):

...if you would be so kind

OpenStudy (zarkon):

if you know f'(x) and that f(0)=0 then finding f(x) is really easy :)

OpenStudy (zarkon):

\[f'(x)=\lim_{y\to x}\frac{f(y)-f(x)}{y-x}\]

OpenStudy (zarkon):

\[f(y)-f(x)=f(y)+f(-x)\]

OpenStudy (zarkon):

does that help?

OpenStudy (zarkon):

ok

OpenStudy (kinggeorge):

If it helps, it's easier for me to see it when I write the derivative as \[f'(x)=\lim_{y\to0} \frac{f(x+y)-f(x)}{y}\]

OpenStudy (turingtest):

that would be \[\lim_{y\to0}{f(y)+x^2y+xy^2\over y}=1+\lim...\]ah... it's \(f'(x)=x^2+1\), right ? :)

OpenStudy (zarkon):

yes

OpenStudy (turingtest):

sweet, thanks y'all I still want to see how your way was gonna go @Zarkon , but I'll work on that myself for a while

OpenStudy (zarkon):

gotta go...fun problem...I'll post it later if you like

OpenStudy (turingtest):

always a pleasure :)